# Non-linear Welfare-Aware Strategic Learning

### v1

@@ Abstract
This paper studies algorithmic decision-making in the presence of strategic individual behaviors, where an ML model is used to make decisions about human agents and the latter can adapt their behavior strategically to improve their future data. Existing results on strategic learning have largely focused on the linear setting where agents with linear labeling function best respond to a (noisy) linear decision policy. Instead, this work focuses on general non-linear settings where agents respond to the decision policy with only "local information" of the policy. Moreover, we simultaneously consider objectives of maximizing decision-maker welfare (model prediction accuracy), social welfare (agent improvement caused by strategic behaviors), and agent welfare (the extent that ML underestimates the agents). We first generalize the agent best response model in previous works to the non-linear setting, then reveal the compatibility of welfare objectives. We show the three welfare can attain the optimum simultaneously only under restrictive conditions which are challenging to achieve in non-linear settings. The theoretical results imply that existing works solely maximizing the welfare of a subset of parties inevitably diminish the welfare of the others. We thus claim the necessity of balancing the welfare of each party in non-linear settings and propose an irreducible optimization algorithm suitable for general strategic learning. Experiments on synthetic and real data validate the proposed algorithm.
@@ 1 Introduction
When machine learning (ML) models are used to automate decisions about human agents (e.g., in lending, hiring, and college admissions), they often have the power to affect agent behavior and reshape the population distribution. With the (partial) information of the decision policy, agents can change their data strategically to receive favorable outcomes, e.g., they may exert efforts to genuinely improve their labels or game the ML models by manipulating their features without changing the labels. This phenomenon has motivated an emerging line of work on Strategic Classification (Hardt et al., 2016), which explicitly takes the agents’ strategic behaviors into account when learning ML models.
@@
However, most existing works on strategic classification solely consider the welfare of one party. For example, Dong et al. (2018); Braverman and Garg (2020); Jagadeesan et al. (2021); Sundaram et al. (2021); Zhang et al. (2022); Eilat et al. (2022) only focus on decision-maker welfare by designing accurate classifiers robust to strategic manipulation. Kleinberg and Raghavan (2020); Harris et al. (2021); Ahmadi et al. (2022a) solely consider maximizing social welfare by designing an optimal policy to incentivize the maximum improvement of agents. To the best of our knowledge, only a few works consider both decision-maker welfare and social welfare (Bechavod et al., 2022; Barsotti et al., 2022). Among them, Bechavod et al. (2022) initiate a theoretical study to understand the relationship between the two welfare, where they identify a sufficient condition under which the decision-maker welfare and social welfare can be optimized at the same time. However, this paper still focuses on incentivizing agent improvement to maximize social welfare.
Moreover, most prior works limit the scope of analysis to linear settings by assuming both the agent labeling function and the ML decision rules are linear Dong et al. (2018); Chen et al. (2020b); Ahmadi et al. (2021); Horowitz and Rosenfeld (2023). This is unrealistic in real applications when the relationship between features and labels is highly intricate and non-linear. Meanwhile, previous works typically modeled strategic behaviors as "(noisy) best responses" to the decision policy Hardt et al. (2016); Jagadeesan et al. (2021); Dong et al. (2018); Chen et al. (2020b); Ahmadi et al. (2021); Horowitz and Rosenfeld (2023), which is not practical when the decision policy is highly complex and agents’ knowledge about the policy (e.g., a neural network) is far from complete. It turns out that the existing theoretical results and methods proposed under linear settings where agents "best respond" can easily be violated in non-linear settings. Thus, there remain gaps and challenges in developing a comprehensive framework for strategic learning under general non-linear settings.
@@
Under general non-linear settings and the proposed agent response model, we first revisit the theoretical results in Bechavod et al. (2022) and show that those results can easily be violated under more complex settings. We then explore the relationships between the objectives of maximizing decision-maker, social, and agent welfare. The results show that the welfare of different parties can only be aligned under rigorous conditions depending on the ground-truth labeling function, agents’ knowledge about the decision policy, and the realizability (i.e., whether the hypothesis class of the decision policy includes the ground-truth labeling function). It implies that existing works that solely maximize the welfare of a subset of parties inevitably diminish the welfare of the others. We thus propose an algorithm to balance the welfare of all parties.
The rest of the paper is organized as follows. After briefly reviewing the related work in Sec. 2, we formulate a general strategic learning setting in Sec. 3, where the agent response and the welfare of the decision-maker, agent, and society are presented. In Sec. 4, we discuss how the existing theoretical results derived under linear settings are violated under non-linear settings and provide a comprehensive analysis of the decision-maker/agent/social welfare; we show all the conditions when they are guaranteed to be compatible in Thm. 4.1. This demonstrates the necessity of an "irreducible" optimization framework balancing the welfare of all parties, which we introduce in Sec. 5. Finally, in Sec. 6, we provide comprehensive experimental results to compare the proposed method with other benchmark algorithms on both synthetic and real datasets.
@@ 2.1 Strategic Learning
Hardt et al. (2016) was the first to formulate strategic classification as a Stackelberg game where the decision-maker publishes a policy first following which agents best respond. A large line of subsequent research has focused on designing accurate classifiers robust to strategic behaviors. For example, Ben-Porat and Tennenholtz (2017); Chen et al. (2020a); Tang et al. (2021) designed robust linear regression predictors for strategic classification, while Dong et al. (2018) proposed an online learning algorithm for the problem. Levanon and Rosenfeld (2021) designed a protocol for the decision maker to learn agent best response when the decision policy is non-linear, but it assumes agents are perfectly aware of the complex classifiers deployed on them. Braverman and Garg (2020); Jagadeesan et al. (2021) relaxed the assumption that the best responses of agents are deterministic and studied how the decision policy is affected by randomness. Sundaram et al. (2021) provided theoretical results on PAC learning under the strategic setting, while Levanon and Rosenfeld (2022); Horowitz and Rosenfeld (2023) proposed a new loss function and algorithms for linear strategic classification. Miller et al. (2020); Shavit et al. (2020); Alon et al. (2020) focused on causal strategic learning where features are divided into "causal" and "non-causal" ones. However, all these works study strategic classification under linear settings and the primary goal is to maximize decision-maker welfare. In contrast, our work considers the welfare of not only the decision-maker but also the agent and society. Another line of work related to strategic learning is Performative Prediction Perdomo et al. (2020) which is a model to capture general model-dependent distribution shifts. Although this framework can model strategic learning settings when the decision policy is strongly convex, it is highly abstract and hard to interpret when considering the welfare of different parties. We refer to Hardt and Mendler-Dünner (2023) as a comprehensive survey for Performative Prediction.
@@ 2.2 Promote Social Welfare in Strategic Learning
Although the definition of social welfare is sometimes subtle Florio (2014), social welfare is often regarded as the improvement of ground truth qualifications of agents in strategic learning settings and we need to consider how decision rules impact social welfare by shaping agent best response. A rich line of works are proposed to design mechanisms to incentivize agents to improveLiu et al. (2019); Rosenfeld et al. (2020); Shavit et al. (2020); Alon et al. (2020); Zhang et al. (2020); Chen et al. (2020a); Kleinberg and Raghavan (2020); Bechavod et al. (2021); Ahmadi et al. (2022a, b); Raab and Liu (2021). The first prominent work was done by Kleinberg and Raghavan (2020), where they studied the conditions under which a linear mechanism exists to incentivize agents to improve their features instead of manipulating them. Specifically, they assume each agent has an effort budget and can choose to invest any amount of effort on either improvement or manipulation, and the mechanism aims to incentivize agents to invest all available efforts in improvement. Harris et al. (2021) designed a more advanced stateful mechanism considering the accumulative effects of improvement behaviors. They stated that improvement can accumulate over time and benefit agents in the long term, while manipulation cannot accumulate state to state. On the contrary, Estornell et al. (2021) considered using an audit mechanism to disincentivize manipulation which instead incentivize improvement. Besides mechanism design, Rosenfeld et al. (2020) studied incentivizing improvement in machine learning algorithms by adding a look-ahead regularization term to favor improvement. Chen et al. (2020a) divided the features into immutable, improvable, and manipulable features and explored linear classifiers that can prevent manipulation and encourage improvement. Jin et al. (2022) also focused on incentivizing improvement and proposed a subsidy mechanism to induce improvement actions and improve social well-being metrics. Bechavod et al. (2021) demonstrated the ability of strategic decision-makers to distinguish features influencing the label of individuals under an online setting. Ahmadi et al. (2022a) proposed a linear model where strategic manipulation and improvement are both present. Barsotti et al. (2022) conducted several empirical experiments when improvement and manipulation are possible and both actions incur a linear deterministic cost.
@@
The works closest to ours are Bechavod et al. (2022); Rosenfeld et al. (2020). Bechavod et al. (2022) provided analytical results of maximizing total improvement without harming anyone’s current qualification, but they assumed both the ground truth and the decision rule are linear. Rosenfeld et al. (2020) aimed to maximize accuracy while punishing the classifier only if it brings the negative externality in non-linear settings. However, solely punishing negative externality is far from ideal. Our work goes further by analyzing the relationships between the welfare of different parties and providing an optimization algorithm to balance the welfare.
@@ 2.3 Welfare-Aware Machine Learning
The early work on welfare-aware machine learning typically considers balancing social welfare and another competing objective such as fairness or profit. For example, Ensign et al. (2018) focused on long-term social welfare when myopically maximizing efficiency in criminal inspection practice creates a runaway feedback loop. Liu et al. (2020); Zhang et al. (2020); Guldogan et al. (2022) showed that instant fairness constraints may harm social welfare in the long run since the agent population may react to the classifier unexpectedly. Rolf et al. (2020) proposed a framework to analyze the Pareto-optimal policy balancing the prediction profit and a possibly competing welfare metric, but it assumed a simplified setting where the welfare of admitting any agent is constant, which is not the case in strategic learning.
@@ 3 Problem Formulation
Consider a population of strategic agents with features $X=(X_{1},...,X_{d})$, $\text{dom}(X)\subset\operatorname{\mathbb{R}}^{d}$, and label $Y\in\{0,1\}$ that indicates agent qualification for certain tasks (with “0" being unqualified and “1" qualified). Denote $P_{X}$ as the probability density function of $X$, and assume labeling function $h(x):=P(Y=1|X=x)$ is continuous. The agents are subject to certain algorithmic decisions $D(x):\mathbb{R}^{d}\to\{0,1\}$ (with “1" being positive and “0" negative), which are made by a decision-maker using a scoring function $f(x):=P(D(x)=1|X=x)$. Unlike many studies in strategic classification that assume $f\in\mathcal{F}$ is a linear function, we allow $\mathcal{F}$ to be non-linear. Moreover, it is possible that hypothesis class $\mathcal{F}$ does not include $h$.
@@ 3.1 Information level and generalized agent response
We consider practical settings where the decisions deployed on agents have downstream effects and can reshape the agent’s future data. We assume agents are honest so that their future labels also change accordingly based on the fixed labeling function $h(x)$. Given a function $f$, we define agent best response to $f$ adhering to previous works on strategic learning Hardt et al. (2016); Dong et al. (2018); Ahmadi et al. (2021) as follows.
@@
It is worth noting that most existing works in strategic learning assume agents either best respond to the exact $f$ (e.g., Hardt et al. (2016); Horowitz and Rosenfeld (2023); Dong et al. (2018); Ahmadi et al. (2021)) or noisy $f$ with additive noise (e.g., Levanon and Rosenfeld (2022); Jagadeesan et al. (2021)). However, the best response model in Def. 3.1 requires the agents to know the exact values of $f(x^{\prime})$ for all $x^{\prime}$ in the feature domain, which is a rather strong requirement and unlikely holds in practice, especially when policy $f\in\mathcal{F}$ is complex (e.g., logistic regression/neural networks). Instead, it may be more reasonable to assume agents only have "local information" about the policy $f$, i.e., agents with features $x$ only have some information about $f$ at $x$ and they know how to improve their qualifications based on their current standings Zhang et al. (2022); Raab and Liu (2021). To formalize this, we define the agent’s information level and the resulting agent response in Def. 3.2.
@@
Agents with features $x$ and information level $K$ know the gradient of $f$ at $x=[x_{1},...,x_{d}]$ up to the $K^{th}$ order (the agents know $\{\nabla f(x),\nabla^{2}f(x),...,\nabla^{K}f(x)\}$ if all these gradients exist). Let $Q_{x}^{f}$ be the Taylor expansion of $f$ at $x$ up to the $K^{th}$ order, i.e.,
$Q_{x}^{f}(x^{\prime})=$ $f(x)+\sum_{i=1}^{d}\frac{\partial f}{\partial x_{i}}(x)(x^{\prime}_{i}-x_{i})+\frac{1}{2!}\sum_{i=1}^{d}\sum_{j=1}^{d}\frac{\partial^{2}f}{\partial x_{i}\partial x_{j}}(x)(x^{\prime}_{i}-x_{i})(x^{\prime}_{j}-x_{j})+\cdots$ (1)
When $f$ is only $k<K$ times differentiable, $Q_{x}^{f}$ simply uses all existing gradient information. Then the agents with information level $K$ will respond to $f$ by first estimating $f$ using $Q_{x}^{f}$, and then best responding to $Q_{x}^{f}$.
@@
It is worth noting that the agent response under $K=1$ is also similar to Rosenfeld et al. (2020) where agents respond to the decision policy by moving toward the direction of the local first-order gradient, and Dean et al. (2023) where the agents aim to reduce their risk based on the available information at the current state. Moreover, Def. 3.2 can also be regarded as a generalization of the previous models where agents best respond to the exact decision policy under the linear setting. To illustrate this, assume the agents have an information level $K\geq 1$ and the linear policy is $f(x)=\beta^{T}x$, then they will estimate $f$ using $Q^{f}_{x}(x^{\prime})=f(x)+\nabla f(x)\cdot(x^{\prime}-x)=\beta^{T}x^{\prime}$, which is the same as $f$. However, for complex $f$, agents with information level $K$ may no longer best respond to $f$. Prop. 3.3 illustrates the necessary and sufficient conditions for an agent to best respond.

### v2

@@ Abstract
This paper studies algorithmic decision-making in the presence of strategic individual behaviors, where an ML model is used to make decisions about human agents and the latter can adapt their behavior strategically to improve their future data. Existing results on strategic learning have largely focused on the linear setting where agents with linear labeling functions best respond to a (noisy) linear decision policy. Instead, this work focuses on general non-linear settings where agents respond to the decision policy with only "local information" of the policy. Moreover, we simultaneously consider objectives of maximizing decision-maker welfare (model prediction accuracy), social welfare (agent improvement caused by strategic behaviors), and agent welfare (the extent that ML underestimates the agents). We first generalize the agent best response model in previous works to the non-linear setting and then investigate the compatibility of welfare objectives. We show the three welfare can attain the optimum simultaneously only under restrictive conditions which are challenging to achieve in non-linear settings. The theoretical results imply that existing works solely maximizing the welfare of a subset of parties usually diminish the welfare of others. We thus claim the necessity of balancing the welfare of each party in non-linear settings and propose an irreducible optimization algorithm suitable for general strategic learning. Experiments on synthetic and real data validate the proposed algorithm.
@@ 1 Introduction
When machine learning (ML) models are used to automate decisions about human agents (e.g., in lending, hiring, and college admissions), they often have the power to affect agent behavior and reshape the population distribution. With the (partial) information of the decision policy, agents can change their data strategically to receive favorable outcomes, e.g., they may exert efforts to genuinely improve their labels or game the ML models by manipulating their features without changing the labels. This phenomenon has motivated an emerging line of work on Strategic Classification (Hardt et al. 2016), which explicitly takes the agents’ strategic behaviors into account when learning ML models.
@@
However, most existing works on strategic classification solely consider the welfare of one party. For example, Dong et al. (2018); Braverman and Garg (2020); Jagadeesan, Mendler-Dünner, and Hardt (2021); Sundaram et al. (2021); Zhang et al. (2022); Eilat et al. (2022) only focus on decision-maker welfare by designing accurate classifiers robust to strategic manipulation. Kleinberg and Raghavan (2020); Harris, Heidari, and Wu (2021); Ahmadi et al. (2022a) solely consider maximizing social welfare by designing an optimal policy to incentivize the maximum improvement of agents. To the best of our knowledge, only a few works consider both decision-maker welfare and social welfare (Bechavod et al. 2022; Barsotti, Kocer, and Santos 2022). Among them, Bechavod et al. (2022) initiate a theoretical study to understand the relationship between the two welfare, where they identify a sufficient condition under which the decision-maker welfare and social welfare can be optimized at the same time. This paper still focuses on incentivizing agent improvement to maximize social welfare.
Moreover, most prior works limit the scope of analysis to linear settings by assuming both the agent labeling function and the ML decision rules are linear (Dong et al. 2018; Chen, Liu, and Podimata 2020; Ahmadi et al. 2021; Horowitz and Rosenfeld 2023). This is unrealistic in real applications when the relationship between features and labels is highly intricate and non-linear. Meanwhile, previous works typically modeled strategic behaviors as "(noisy) best responses" to the decision policy (Hardt et al. 2016; Jagadeesan, Mendler-Dünner, and Hardt 2021; Dong et al. 2018; Chen, Liu, and Podimata 2020; Ahmadi et al. 2021; Horowitz and Rosenfeld 2023), which is not practical when the decision policy is highly complex and agents’ knowledge about the policy (e.g., a neural network) is far from complete. It turns out that the existing theoretical results and methods proposed under linear settings where agents "best respond" can easily be violated in non-linear settings. Thus, there remain gaps and challenges in developing a comprehensive framework for strategic learning under general non-linear settings.
@@
Under general non-linear settings and the proposed agent response model, we first revisit the theoretical results in Bechavod et al. (2022) and show that those results can easily be violated under more complex settings. We then explore the relationships between the objectives of maximizing decision-maker, social, and agent welfare. The results show that the welfare of different parties can only be aligned under rigorous conditions depending on the ground-truth labeling function, agents’ knowledge about the decision policy, and the realizability (i.e., whether the hypothesis class of the decision policy includes the ground-truth labeling function). It implies that existing works that solely maximize the welfare of a subset of parties may inevitably diminish the welfare of the others. We thus propose an algorithm to balance the welfare of all parties.
The rest of the paper is organized as follows. After reviewing the related work in Sec. 2, we formulate a general strategic learning setting in Sec. 3, where the agent response and the welfare of the decision-maker, agent, and society are presented. In Sec. 4, we discuss how the existing theoretical results derived under linear settings are violated under non-linear settings and provide a comprehensive analysis of the decision-maker/agent/social welfare; we show all the conditions when they are guaranteed to be compatible in Thm. 4.1. This demonstrates the necessity of an "irreducible" optimization framework balancing the welfare of all parties, which we introduce in Sec. 5. Finally, in Sec. 6, we provide comprehensive experimental results to compare the proposed method with other benchmark algorithms on both synthetic and real datasets.
@@ 2.1 Strategic Learning
Hardt et al. (2016) was the first to formulate strategic classification as a Stackelberg game where the decision-maker publishes a policy first following which agents best respond. A large line of subsequent research has focused on designing accurate classifiers robust to strategic behaviors. For example, Ben-Porat and Tennenholtz (2017); Chen, Wang, and Liu (2020); Tang, Ho, and Liu (2021) designed robust linear regression predictors for strategic classification, while Dong et al. (2018); Ahmadi et al. (2021); Lechner, Urner, and Ben-David (2023); Shao, Blum, and Montasser (2024) studied online learning algorithms for the problem. Levanon and Rosenfeld (2021) designed a protocol for the decision maker to learn agent best response when the decision policy is non-linear, but it assumes agents are perfectly aware of the complex classifiers deployed on them. Braverman and Garg (2020); Jagadeesan, Mendler-Dünner, and Hardt (2021) relaxed the assumption that the best responses of agents are deterministic and studied how the decision policy is affected by randomness. Sundaram et al. (2021) provided theoretical results on PAC learning under the strategic setting, while Levanon and Rosenfeld (2022); Horowitz and Rosenfeld (2023) proposed a new loss function and algorithms for linear strategic classification. Miller, Milli, and Hardt (2020); Shavit, Edelman, and Axelrod (2020); Alon et al. (2020) focused on causal strategic learning where features are divided into "causal" and "non-causal" ones. However, all these works study strategic classification under linear settings and the primary goal is to maximize decision-maker welfare. In contrast, our work considers the welfare of not only the decision-maker but also the agent and society. Another line of work related to strategic learning is Performative Prediction (Perdomo et al. 2020; Jin et al. 2024) to capture general model-dependent distribution shifts. Despite the generality, the framework is highly abstract and hard to interpret when considering the welfare of different parties. We refer to Hardt and Mendler-Dünner (2023) as a comprehensive survey.
@@ 2.2 Promote Social Welfare in Strategic Learning
Although the definition of social welfare is sometimes subtle (Florio 2014), social welfare is often regarded as the improvement of ground truth qualifications of agents in strategic learning settings and we need to consider how decision rules impact social welfare by shaping agent best response. A rich line of works are proposed to design mechanisms to incentivize agents to improve (e.g., Liu et al. (2019); Shavit, Edelman, and Axelrod (2020); Alon et al. (2020); Zhang et al. (2020); Chen, Wang, and Liu (2020); Kleinberg and Raghavan (2020); Ahmadi et al. (2022a, b); Jin et al. (2022); Xie and Zhang (2024)). The first prominent work was done by Kleinberg and Raghavan (2020), where they studied the conditions under which a linear mechanism exists to incentivize agents to improve their features instead of manipulating them. Specifically, they assume each agent has an effort budget and can choose to invest any amount of effort on either improvement or manipulation, and the mechanism aims to incentivize agents to invest all available efforts in improvement. Harris, Heidari, and Wu (2021) designed a more advanced stateful mechanism considering the accumulative effects of improvement behaviors. They stated that improvement can accumulate over time and benefit agents in the long term, while manipulation cannot accumulate state to state. On the contrary, Estornell, Das, and Vorobeychik (2021) considered using an audit mechanism to disincentivize manipulation which instead incentivize improvement. Besides mechanism design, Rosenfeld et al. (2020) studied incentivizing improvement in machine learning algorithms by adding a look-ahead regularization term to favor improvement. Chen, Wang, and Liu (2020) divided the features into immutable, improvable, and manipulable features and explored linear classifiers that can prevent manipulation and encourage improvement. Jin et al. (2022) also focused on incentivizing improvement and proposed a subsidy mechanism to induce improvement actions and improve social well-being metrics. Bechavod et al. (2021) demonstrated the ability of strategic decision-makers to distinguish features influencing the label of individuals under an online setting. Ahmadi et al. (2022a) proposed a linear model where strategic manipulation and improvement are both present. Barsotti, Kocer, and Santos (2022) conducted several empirical experiments when improvement and manipulation are possible and both actions incur a linear deterministic cost.
@@
The works closest to ours are Bechavod et al. (2022); Rosenfeld et al. (2020). Bechavod et al. (2022) provided analytical results of maximizing total improvement without harming anyone’s current qualification, but they assumed both the ground truth and the decision rule are linear. Rosenfeld et al. (2020) maximized accuracy while punishing the classifier only if it brings the negative externality in non-linear settings. However, solely punishing negative externality is far from ideal. Our work goes further by analyzing the relationships between the welfare of different parties and providing an optimization algorithm to balance the welfare.
@@ 2.3 Welfare-Aware Machine Learning
The early work on welfare-aware machine learning typically considers balancing social welfare and another competing objective such as fairness or profit. For example, Ensign et al. (2018) focused on long-term social welfare when myopically maximizing efficiency in criminal inspection practice creates a runaway feedback loop. Liu et al. (2020); Zhang et al. (2020); Guldogan et al. (2022) showed that instant fairness constraints may harm social welfare in the long run since the agent population may react to the classifier unexpectedly. Rolf et al. (2020) proposed a framework to analyze the Pareto-optimal policy balancing the prediction profit and a possibly competing welfare metric but assumed a simplified setting where the welfare of admitting any agent is constant which is not the case in strategic learning.
@@ 3 Problem Formulation
Consider a population of strategic agents with features $X=(X_{1},...,X_{d})$, $\text{dom}(X)\subset\operatorname{\mathbb{R}}^{d}$, and label $Y\in\{0,1\}$ that indicates agent qualification for certain tasks (with “0" being unqualified and “1" qualified). Denote $P_{X}$ as the probability density function of $X$, and assume labeling function $h(x):=P(Y=1|X=x)$ is continuous. The agents are subject to certain algorithmic decisions $D(x):\mathbb{R}^{d}\to\{0,1\}$ (with “1" being positive and “0" negative), which are made by a decision-maker using a scoring function $f(x):=P(D(x)=1|X=x)$. Unlike many studies in strategic classification that assume $f\in\mathcal{F}$ is a linear function [^1] A linear scoring function can be $\mathbf{1}(w^{T}x+b\geq 0)$ (e.g., Levanon and Rosenfeld (2022); Horowitz and Rosenfeld (2023)) or $w^{T}x+b$ when $\text{dom}(X)$ is bounded (e.g., Bechavod et al. (2022))., we allow $\mathcal{F}$ to be non-linear. Moreover, it is possible that hypothesis class $\mathcal{F}$ does not include $h$.
@@ 3.1 Information level and generalized agent response
We consider practical settings where the decisions deployed on agents have downstream effects and can reshape the agent’s future data. We also focus on "benign" agents whose future labels change according to the fixed labeling function $h(x)$ (Raab and Liu 2021; Rosenfeld et al. 2020; Guldogan et al. 2022). Given a function $f$, we define agent best response to $f$ in Def. 3.1, which is consistent with previous works on strategic learning (Hardt et al. 2016; Dong et al. 2018; Ahmadi et al. 2021).
@@
It is worth noting that most existing works in strategic learning assume agents either best respond to the exact $f$ (e.g., Hardt et al. (2016); Horowitz and Rosenfeld (2023); Dong et al. (2018); Ahmadi et al. (2021)) or noisy $f$ with additive noise (e.g., Levanon and Rosenfeld (2022); Jagadeesan, Mendler-Dünner, and Hardt (2021)). However, the best response model in Def. 3.1 requires the agents to know the exact values of $f(x^{\prime})$ for all $x^{\prime}$ in the feature domain, which is a rather strong requirement and unlikely holds in practice, especially when policy $f\in\mathcal{F}$ is complex (e.g., logistic regression/neural networks). Instead, it may be more reasonable to assume agents only have "local information" about the policy $f$, i.e., agents with features $x$ only have some information about $f$ at $x$ and they know how to improve their qualifications based on their current standings (Zhang et al. 2022; Raab and Liu 2021). To formalize this, we define the agent’s information level and the resulting agent response in Def. 3.2.
@@
Agents with features $x$ and information level $K$ know the gradient of $f$ at $x=[x_{1},...,x_{d}]$ up to the $K^{th}$ order (the agents know $\{\nabla f(x),\nabla^{2}f(x),...,\nabla^{K}f(x)\}$ if all these gradients exist). Let $Q_{x}^{f}$ be the Taylor expansion of $f$ at $x$ up to the $K^{th}$ order. When $f$ is only $k<K$ times differentiable, $Q_{x}^{f}$ simply uses all existing gradient information. Then the agents with information level $K$ will respond to $f$ by first estimating $f$ using $Q_{x}^{f}$, and then best responding to $Q_{x}^{f}$.
@@
It is worth noting that the agent response under $K=1$ is also similar to Rosenfeld et al. (2020) where agents respond to the decision policy by moving toward the direction of the local first-order gradient, and Dean et al. (2023) where the agents aim to reduce their risk based on the available information at the current state. Moreover, Def. 3.2 can also be regarded as a generalization of the previous models where agents best respond to the exact decision policy under the linear setting. To illustrate this, assume the agents have $K\geq 1$ and the linear policy is $f(x)=\beta^{T}x$, then the agents with feature $x$ will estimate $f$ using $Q^{f}_{x}(x^{\prime})=f(x)+\nabla f(x)\cdot(x^{\prime}-x)=\beta^{T}x^{\prime}$, which is the same as $f$. However, for complex $f$, agents with information level $K$ may no longer best respond to $f$. Prop. 3.3 illustrates the necessary and sufficient conditions for an agent to best respond to $f$.